OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 2 -1 1 -2

OpenStudy (anonymous):

mulitply first by 4 and second by -3

OpenStudy (anonymous):

2

OpenStudy (anonymous):

sorry second by 3

hero (hero):

Add both equations together: 5x - y = 11 y = 5x - 11 substitute that into the first equation 2x + 3(5x - 11) = 1 2x + 15x - 33 = 1 17x = 34 x = 34/17 x = 2

OpenStudy (anonymous):

what about this one? For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10 -2y + 10 2y + 10 2y - 10 -2y - 10